# A Survey of Text Watermarking in the Era of Large Language Models

### v1

@@ Abstract.
This survey aims to comprehensively summarize current text watermarking technologies, covering three main aspects: (1) an overview and comparison of different text watermarking techniques; (2) evaluation methods for text watermarking algorithms, including their success rate, impact on text quality, robustness, and unforgeability; (3) potential applications of text watermarking technologys. This survey aims to help researchers thoroughly understanding the text watermarking technologies, thereby fostering further development.
@@ 5.2.3. Task-specified Evaluation
Although the assessment method based on semantic scores could effectively evaluate whether adding watermarks alters the text semantics, its impact on real-world applications remains unclear. Consequently, many studies are now focusing on exploring the effects of watermarking algorithms on specific downstream tasks to assess their impact on text quality. These specific downstream tasks include machine translation (Topkara et al., 2006a; Hu et al., 2023; Wu et al., 2023; Zhao et al., 2023b), sentiment classification (Yang et al., 2023), code generation (Lee et al., 2023; Sun et al., 2023), text summarization (Hu et al., 2023; Wu et al., 2023), and story generation (Zhao et al., 2023b).
@@
Machine Translation. Typically, only watermarking Large Language Models (Section 4) incorporate machine translation as a downstream task for testing text quality. Specifically, the evaluation involves comparing the translation outcomes between a watermarked LLM and the original unwatermarked LLM. This comparison is conducted using the BLEU score, a widely used metric in machine translation. For the choice of translation LLMs, (Hu et al., 2023; Wu et al., 2023) employed the Multilingual BART (Liu et al., 2020), while Takezawa et al. (2023) utilized the NLLB-200 model (Costa-jussà et al., 2022). The translation data typically employed is the WMT14 dataset, where the translations between French, German, and English are the most commonly utilized settings. Most watermarking approaches for LLMs tend to result in a slight decrease in BLEU scores (Kirchenbauer et al., 2023a; Takezawa et al., 2023; Liu et al., 2023b). However, the unbiased watermark methods for LLM (Hu et al., 2023; Wu et al., 2023) exhibit almost no decline in BLEU values, demonstrating the superiority of unbiased watermarking.
@@
Sentiment Classification. Using sentiment classification as a downstream task can validate whether text watermarking algorithms can affect the sentiment distribution, i.e., whether the text can maintain its original sentiment (e.g., positive or negative) after the insertion of a watermark. Yang et al. (2023) analyzed the sentiment distribution of texts with and without watermarks using the Twitter-XLM-RoBERTa-Base-Sentiment  [^2] https://huggingface.co/cardiffnlp/twitter-xlm-roberta-base-sentiment model. Since different sentiments usually have clear distinctions, maintaining sentiment distribution is typically an easy straightforward task for watermark algorithms.
@@
Code Generation. Text watermarking for code generation is an important application, which could test the impact of watermarking on code functionality. Code evaluation could use unit test metrics like pass@k (Kulal et al., 2019), or matching metrics such as BLEU and Exact match. Sun et al. (2023) inserted watermark features into a subset of a dataset to achieve code dataset watermarking. They showed negligible differences in BLEU and exact match scores between models trained on datasets with and without watermarks. However, embedding watermarks into individual code is more challenging than watermarking an entire dataset. Lee et al. (2023) added watermarks to large models for code generation, leading to a nearly 10% performance decline in pass@100 and pass@80 metrics. Given that even minor modifications can disrupt code functionality, and considering code’s inherently low entropy, the code generation task presents a highly challenging downstream task for text quality test.
@@
Story Generation. Similar to text summarization tasks, story generation also presents a suitable scenario for evaluating watermark algorithms for Large Language Models. Tests in the story generation context typically involve inputting the first half of a story and having the model (LLM) predict its ending. The ROCstories dataset (Mostafazadeh et al., 2016) is commonly used, with ROUGE as the evaluation metric. According to the experiments by Zhao et al. (2023b), current watermark algorithms still cause a 1%-2% decrease in performance on ROUGE scores. Moreover, there has been limited work evaluating performance on story generation, suggesting that further testing in this area could be a future direction.

### v2

@@ Abstract.
This survey aims to comprehensively summarize current text watermarking technologies, covering three main aspects: (1) an overview and comparison of different text watermarking techniques; (2) evaluation methods for text watermarking algorithms, including their success rate, impact on text quality, robustness, and unforgeability; (3) potential applications of text watermarking technologies. This survey aims to help researchers thoroughly understanding the text watermarking technologies, thereby fostering further development.
@@ 5.2.3. Task-specified Evaluation
Although the assessment method based on semantic scores could effectively evaluate whether adding watermarks alters the text semantics, its impact on real-world applications remains unclear. Consequently, many studies are now focusing on exploring the effects of watermarking algorithms on specific downstream tasks to assess their impact on text quality. These specific downstream tasks include machine translation (Topkara et al., 2006a; Hu et al., 2023; Wu et al., 2023; Zhao et al., 2023b), sentiment classification (Yang et al., 2023), knowledge understanding (Tu et al., 2023), code generation (Lee et al., 2023; Sun et al., 2023; Tu et al., 2023), text summarization (Hu et al., 2023; Wu et al., 2023; Tu et al., 2023), story generation (Zhao et al., 2023b), question answering (Tu et al., 2023) and instruction following (Tu et al., 2023).
@@
Machine Translation. Typically, only watermarking Large Language Models (Section 4) incorporate machine translation as a downstream task for testing text quality. Specifically, the evaluation involves comparing the translation outcomes between a watermarked LLM and the original unwatermarked LLM. This comparison is conducted using the BLEU score, a widely used metric in machine translation. For the choice of translation LLMs, (Hu et al., 2023; Wu et al., 2023) employed the Multilingual BART (Liu et al., 2020), while Takezawa et al. (2023) utilized the NLLB-200 model (Costa-jussà et al., 2022). The translation data typically employed is the WMT14 dataset, where the translations between French, German, and English are the most commonly utilized settings. Most watermarking approaches for LLMs result in a slight decrease in BLEU scores (Kirchenbauer et al., 2023a; Takezawa et al., 2023; Liu et al., 2023b). However, the unbiased watermark methods (Hu et al., 2023; Wu et al., 2023) exhibit almost no decline in BLEU values, demonstrating the superiority of unbiased watermarking.
@@
Sentiment Classification. Using sentiment classification as a downstream task can validate whether text watermarking algorithms can affect the sentiment distribution, i.e., whether the text can maintain its original sentiment (e.g., positive or negative) after the insertion of a watermark. Yang et al. (2023) analyzed the sentiment distribution of texts with and without watermarks using the Twitter-XLM-RoBERTa-Base-Sentiment  [^2] https://huggingface.co/cardiffnlp/twitter-xlm-roberta-base-sentiment model. Different sentiments generally have clear differences, making it easy for watermark algorithms to maintain sentiment distribution.
Knowledge Understanding. To explore the performance of the watermark for LLMs algorithm in tasks with shorter output lengths, Tu et al. (2023) proposed testing on Knowledge Understanding tasks. Specifically, this involves two scenarios: Knowledge Probing, using the KoLA dataset (Yu et al., 2023) for assessing factual recall in LLMs, and Concept Probing, employing the Copen dataset (Peng et al., 2022) for evaluating conceptual understanding. The typical evaluation metric for these tasks is the F1 score. In practical tests, applying watermarks to Knowledge Understanding tasks significantly decreases the F1 scores across all algorithms, indicating the challenging nature of this scenario.
@@
Code Generation. Text watermarking for code generation is an important application, which could test the impact of watermarking on code functionality. Code evaluation could use unit test metrics like pass@k (Kulal et al., 2019), or matching metrics such as BLEU and Exact match. Sun et al. (2023) inserted watermark features into a subset of a dataset to achieve code dataset watermarking. They showed negligible differences in BLEU and exact match scores between models trained on datasets with and without watermarks. However, embedding watermarks into individual code is more challenging than watermarking an entire dataset. Lee et al. (2023) added watermarks to large models for code generation, leading to a nearly 10% performance decline in pass@100 and pass@80 metrics. Given that even minor modifications can disrupt code functionality, and considering code’s inherently low entropy, the code generation task presents a highly challenging downstream task for text quality test (Tu et al., 2023).
@@
Story Generation. Similar to text summarization tasks, story generation also presents a suitable scenario for evaluating watermark algorithms for Large Language Models. Tests in the story generation context typically involve inputting the first half of a story and having the model (LLM) predict its ending. The ROCstories dataset (Mostafazadeh et al., 2016) is commonly used, with ROUGE as the evaluation metric. According to the experiments by Zhao et al. (2023b), current watermark algorithms still cause a 1%-2% decrease in performance on ROUGE scores. Furthermore, minimal research has been conducted on story generation performance, indicating potential for future exploration.
Question Answering. Question answering (QA) is an important downstream application of Large Language Models, and testing watermark algorithms for LLMs on this task is equally crucial. Tu et al. (2023) conducted tests on three different QA tasks, specifically: the ELI5 dataset (Fan et al., 2019) for long-form QA tasks, the FinQA dataset (Maia et al., 2018) for Finance QA tasks, and the HotpotQA dataset (Yang et al., 2018) for multi-hop reasoning QA tasks. For the long-form QA and Finance QA tasks, the Rouge-L metric was used for evaluation, while the F1 score was utilized for the multi-hop reasoning QA task. Experimental results revealed that, after the introduction of watermarks, all current watermark algorithms experienced a performance decline of about 50% in Question Answering tasks, indicating the challenging nature of watermark algorithms in the context of Question Answering.
Instruction Following. The instruction following ability of Large Language Models has recently become an especially important aspect to assess, reflecting whether LLMs can accurately follow user instructions to generate output in open-ended situations. Tu et al. (2023) tested the impact of the current watermark for LLMs algorithm on the Instruction Following task using the AlpacaFarm dataset (Dubois et al., 2023). The evaluation method adopted was GPT4-Judge (Zheng et al., 2023), where the GPT-4 model judges which output, between the watermarked LLM and Davinci-003, is better in response to a given instruction. Under this metric, both Llama2-7B-chat and Internlm-7B-8k models showed over a 90% decline in performance with watermark, indicating that instruction following presents a particularly challenging scenario for watermark algorithms.